OpenStudy (anonymous):

A substance with a hydroxide ion concentration of 1.55 x 10-11 M will have a pH of ___. I got 3.19 is that right?

OpenStudy (frostbite):

2 sec need to get my hands on a cal.

OpenStudy (anonymous):

ok please cuz I have like 4 more like this lol

OpenStudy (anonymous):

it's just I want to be sure I did them right, Thanks so much!

OpenStudy (frostbite):

Yea it is right. :)

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

my next one would be What is the hydronium ion concentration of a solution with a pH of 2.57? I got 2.7 * 10^-3

OpenStudy (frostbite):

Correct :)

OpenStudy (frostbite):

Any more? :)

OpenStudy (anonymous):

yup but just a quick question, Molar mass is the number of grams of a given compound in one mole of the compound and 6.02 x 10^23 particles of the compound. Or is it just one mole of a compound?

OpenStudy (anonymous):

@Frostbite u there?

OpenStudy (frostbite):

Yea I'm here. just doing multiple things are the same time. Molar mass is the number of grams for 1 mol of the compound yes. Also what the definition say:\[\Large M=\frac{ m }{ n }\] Set \(n=1\)\[\Large M=m\] And when we say we got 1 mole of molecules we then mean we got 6.02 x 10^23 molecules yes.

OpenStudy (anonymous):

ok my next question is You have 350 mL of a gas at a pressure of 2.4 atm. If the volume is increased to 575 mL, what is the new pressure? I got .68 atm is that right?

OpenStudy (frostbite):

Emmm. \[\Large p _{1}V _{1}=p _{2}V _{2}\] Did you use that equation?

OpenStudy (anonymous):

yes but i used my calculator

OpenStudy (anonymous):

nvm got it :)

OpenStudy (frostbite):

Hehe well I can only do it in my head: \[\Large p _{2}=\frac{ p _{1}V _{1} }{ V _{2} }=\frac{ 350 ml \times 2.4 atm }{ 575ml }\]

OpenStudy (frostbite):

it was right?

OpenStudy (anonymous):

nope

OpenStudy (frostbite):

1.46 atm instead?

OpenStudy (anonymous):

yup

OpenStudy (frostbite):

The head (almost) never fails ;)

OpenStudy (frostbite):

Anything else I can assist with? :)